AS8096(1L) AS8096(1L) NAME as8096,as80196−crossassemblersformicrocomputers SYNOPSIS as8096[-pcpu][-llistfile][-ohexfile][-d][-ssymbolfile]input as80196[-pcpu][-llistfile][-ohexfile][-d][-ssymbolfile]input DESCRIPTION Theas8096,as80196commandsassemblestheinputfileintoatextoutputfilerepresentingthe programmemoryforamicrocomputer. Options -pcpuOverridetheinstructionsetselection. ValidValuesforCpu 8096,80196 -llistfile Outputafileformatedwiththememoryaddress,data,andsourceinputlines. -ohexfile Outputthememorydataimageinaformacceptedbymostpromprogrammers. -hhexfile Sameas−o. −dSavetheintermediatedatafile(seeFILE)andaborttheexecutionattheendofprocessing. -ssymbolfile Printthesymboltablevaluesandnames,oneperlineinthespecifiedfile. FILES /usr/tmp/frtXXXXXX SEEALSO FrankensteinCrossAssemblersUsersManual(filebase.doc)Appendixforas8096Frankenstein Assembler(fileas8096.doc) NOTES Thereisonlyoneinputfile. Theprogramcanselectwhichsubsetofinstructionsisaccepted.Theprogramfirstusesthename withwhichitisinvoked.Thisisscannedforasubstringwhichindicateswhichsettouse.The-p optionsoverridesthisselctionbyperformingthesamesubstringsearch.Finallytheinputfilecan selectwhichsubsettousewiththeCPUstatement. Thereshouldonlybeoneexecutablefilewithallofitsnameslinkedtoit. 1